Two charged spheres placed 43 cm apart exert a force of 1.40 × 10-14 N on each other. If one of the spheres has a charge of 1.68 × 10-17 C, what is the charge of the other sphere?




1
Expert's answer
2021-05-09T13:06:48-0400

The force between two point charges obeys the Coulomb's law:


"F = k\dfrac{|q_1||q_2|}{r^2}"

where "q_1 = 1.68\times10^{-17}C,\space q_2" are the charges of the first and second spheres respectively, "r = 43cm = 0.43m" is the distance between the spheres, and "k = 9\times 10^9 N\cdot m^2/C^2" is the Coulomb's constant. Expressing the second charge and substituting "F = 1.4\times 10^{-14}N" , obtain:


"q_2 = \dfrac{Fr^2}{kq_1}\\
q_2 = \dfrac{1.4\times 10^{-14}\cdot 0.43^2}{9\times 10^9\cdot 1.68\times 10^{-17}} \approx 1.7\times 10^{-10}C"

Answer. "1.7\times 10^{-10}C".